Two blocks of mass m = 10.8 kg each are fastened t

Two blocks of mass m = 10.8 kg each are fastened to the ceiling of an elevator.

The elevator accelerates upward at a = 2.00 m/s2. Find the tension in the bottom rope and the top rope.
Answer #1

assume the masses are hung in series (one below the other).
T = m(a + g), so for the top rope,
T = 2 * 10.8 kg * (9.8 + 2 )m/s2 = 255 N